# MDL: A Unified Multi-Distribution Learner in Large-scale Industrial Recommendation through Tokenization

### v1

@@ 1. Introduction
To address the above issues, we further draw inspiration from LLMs to enhance MSL and MTL in recommendations. LLMs accomplish specific downstream tasks by inputting appropriate prompt information (OpenAI, 2023; Team, 2025). As a special type of token, prompt tokens can fully activate the vast latent knowledge of LLMs through the self-attention mechanism. Similarly, we posit that the scenario and task information can serve as prompts within recommender models like that in LLMs. By tokenizing scenario and task priors into a unified latent space, we transform them from mere auxiliary inputs into active participants within the core feature-interaction backbone. Instead of restricting scenario/task information to the ”shallows” of the model (e.g., gates or heads), tokenized priors can penetrate deep and multi-layered interaction modules. This allows distribution-specific signals to adaptively modulate the model’s massive parameter space, fully harvesting the benefits of the scaling law (corresponding to the first issue). By projecting both input-side scenario contexts and output-side task objectives into a consistent tokenized format, we bridge the gap between MSL and MTL. This enables a holistic, joint modeling framework where scenario-task dependencies are captured through a unified interaction protocol (corresponding to the second issue).
@@ 3.2.2. Feature-Scenario/Task Token Interaction
The interaction between scenario/task tokens and feature tokens primarily aims to enable different scenario tokens or task tokens to utilize distinct feature information, thereby achieving differentiated modeling for multi-scenario and multi-task settings. Since task and scenario information are tokenized from the bottom layer, they can fully interact with feature information layer by layer and from the bottom up. This approach alleviates the challenge of insufficient utilization of complex feature interaction modules, which was previously limited to the certain intermediate or output layers.
Specifically, we define a type of interaction called Domain-aware Attention between scenario/task token and feature token. Without loss of generality, we take the interaction between task tokens $\bm{T}_{t}$ and feature tokens $\bm{T}_{f}$ as an example to illustrate this module. The Domain-aware Attention can be viewed as a variant of cross-attention. Given the task token $\bm{T}_{t}$, we use it as query, while treating the feature token $\bm{T}_{f}$ as key and value. Then they are interacted through cross multi-head attention as follows:
@@ 3.2.3. Scenario-Task Token Interaction
Specifically, the Domain-fused Module consists of two steps: (1) instance-level scenario token selection, and (2) hybrid information fusion between task and scenario token. Firstly, for each instance in the dataset, we select the corresponding scenario tokens $\{{\bm{t}_{s,i},\bm{t}_{s,j}}\}$ from the full scenario token set $\bm{T}_{s}=\{\bm{t}_{s,1},\bm{t}_{s,2},\cdots,\bm{t}_{s,{n_{s}}},\bm{t}_{s,global}\}$ based on the scenario information associated with each instance, along with the global scenario token $\bm{t}_{s,global}$. For example, suppose a model is designed to serve $N_{s}$ scenarios. For a given instance, it must belong to one or more of these scenarios (in cases where scenario overlap exists). Based on this information, the corresponding scenario tokens are selected. After selecting the scenario tokens $\{{\bm{t}_{s,i},\bm{t}_{s,j}},\bm{t}_{s,global}\}$, we aggregate their information using a mean pooling approach:
@@ 3.3. MDL Block
Forward Propagation of Scenario Tokens. The information propagation of scenario tokens primarily involves domain-aware attention between scenario token and feature token (Eq 8), per-scenario FFN, and residual connection. It can be formulated as follows:
@@
Forward Propagation of Task Tokens. The information propagation of task tokens primarily consists of domain-aware attention between task token and feature token (Eq 7), domain-fused Module (Eq 9, Eq 10), per-task FFN and residual connection. It can be formulated as follows:
@@ 4.3. Ablation Study
The experimental results of the MDL variants are reported in Table 2. From it, we can observe that removing each of the components leads to the performance decrease. The variant $w/o$ task token and $w/o$ scenario token indicate the importance of tokenization of scenario information and task information. The variants $w/o$ global scenario token shows the importance of the global token. The variants $w/o$ task-feautre interaction and $w/o$ scenario-feature interaction indicate the effectiveness of the elaborated interaction mechanisms among these tokens.
@@ 4.5. Online Experiments
To further verify the effectiveness of the proposed approach MDL, we conducted an online A/B test on Douyin Search for one month. The core online metrics include Change Query Rate and LT30, as introduction in Section 4.1.2. Table 3 reports the relative improvements in the different scenarios and the overall performance. The proposed method achieved an overall improvement of -0.3267% in change query rate and +0.0626% in LT30. As for each scenario, it also achieve significant improvements on the core metrics. Following the successful experimental results, the proposed approach MDL has been fully deployed on Douyin Search.
@@ 5. Related Work
Large-scale Recommendations. Inspired by the success of scaling laws in large language models (LLMs) (Hoffmann et al., 2022; Kaplan et al., 2020), the scaling laws of recommendation models have recently been widely studied and have achieved significant success in real-world industrial recommendation systems (Zhang et al., 2024a; Zhu et al., 2025; Yan et al., 2025; Zhai et al., 2024; Han et al., 2025a; Zhang et al., 2025; Huang et al., 2026). Typically, these methods increase the model parameter size and computational complexity by stacking various efficient feature interaction structures, thereby improving the recommendation performance. One type of method focuses on scaling the parameters and complexity of the non-sequential feature interaction module for better performance. For example, Wukong (Zhang et al., 2024a) scales the feature interaction module by introducing a stacked factorization machine-based architecture. Rankmixer (Zhu et al., 2025) introduces multi-head token mixing and pertoken feed-forward networks to scale recommender models, while maintaining inference latency and achieving superior performance. FAT (Yan et al., 2025) scales recommendation models by introducing field-aware transformer with field-based interaction priors in attention. Another typical method adopts the transformer-based structure for modeling non-sequential and sequential feature in a unified way and scaling constantly. For instance, HSTU (Zhai et al., 2024) sequentializes the heterogeneous feature space and then performs sequence transduction conditioned on contextual and candidate signals. MTGR (Han et al., 2025a) further scales recommendation models based on HSTU for acceleration and leakage avoidance. OneTrans (Zhang et al., 2025) introduces a unified transformer-based backbone that simultaneously models user-behavior sequences and feature interactions via a unified tokenizer, enabling efficient parameter scaling and superior performance. Compared to these methods, our work further explores how large-scale recommendation models can better perform on multi-scenario and multi-task settings.
@@
Multi-scenario and Multi-task Recommendations. In real-world industrial recommendation systems, a recommendation model usually needs to serve multiple scenarios and tasks. Multi-scenario learning (MSL) and multi-task learning (MTL) in recommender system have been widely studied in previous works (Caruana, 1998; Ma et al., 2018; Misra et al., 2016; Ma et al., 2019; Tang et al., 2020; Li et al., 2020; Sheng et al., 2021; Chang et al., 2023; Zhou et al., 2023; Song et al., 2024; Yan et al., 2024; Ding et al., 2021). For MTL, most existing methods adopt a shared-specific architecture in which task-shared and task-specific components are combined together (e.g., Sharedbottom (Caruana, 1998), Cross-stitch (Misra et al., 2016), SNR (Ma et al., 2019)). In addition, MoE (Mixture-of-Experts) based structures have also been widely proposed, aiming to enhance the effectiveness of multi-task learning by designing various expert mechanisms (e.g., MMoE (Ma et al., 2018), PLE (Tang et al., 2020)). For MSL, it aims to use a unified model to learn data from multiple scenarios, thereby improving recommendation performance. Most existing work also adopts a shared-specific architecture in which scenario-shared and scenario-specific components are combined together. The framework used in MTL can also be applied in MSL by treating the predictions for different scenarios as distinct prediction tasks (e.g., SharedBottom (Caruana, 1998), MMoE (Ma et al., 2018)). In addition, some specialized multi-scenario methods have also been proposed (Li et al., 2020; Sheng et al., 2021; Chang et al., 2023; Zhou et al., 2023; Song et al., 2024; Mu et al., 2023; Zhang et al., 2022a; Yang et al., 2022; Zhang et al., 2022b; Niu et al., 2021; Shen et al., 2021; Chen et al., 2020). For example, HMoE (Li et al., 2020) proposes the hybrid of implicit and explicit MoE structure which learns the scenario and task relationships. STAR (Sheng et al., 2021) introduces a start topology structure where a centered network is used for all scenarios and scenario-specific network is generated for each scenario. PEPNet (Chang et al., 2023) takes the scenario related prior information to adjust the feature embeddings and hidden units through gate mechanisms. In contrast to these approaches which are based on the shared-specific architecture, we propose a novel multi-distribution learning framework from the perspective of tokenization, aiming to enhance the learning capability and unified modeling ability of large-scale recommendation models on multi-scenario and multi-task settings.
@@ 6. Conclusion
In this paper, we propose a novel multi-distribution learning framework for multi-scenario learning (MSL) and multi-task learning (MTL) in large-scale industrial recommender system, named MDL. Different from existing MSL and MTL recommendation methods that are based on a shared-specific architecture, the proposed approach inspired by large language models (LLMs) to leverage prompt-based tokenization to achieve model predictions on various scenarios and tasks settings. Specifically, MDL tokenizes feature, scenario, and task information into a unified tokenized format, and elobrate three key interaction mechanisms among these tokens. Extensive experiments on both offline evaluation and online test have demonstrated the effectiveness. The proposed approach MDL has been fully deployed on Douyin Search platform.

### v2

@@ 1. Introduction
To address the above issues, we further draw inspiration from LLMs to enhance MSL and MTL in recommendations. LLMs accomplish specific downstream tasks by inputting appropriate prompt information (OpenAI, 2023; Team, 2025). As a special type of token, prompt tokens can fully activate the vast latent knowledge of LLMs through the self-attention mechanism. Similarly, we posit that the scenario and task information can serve as prompts within recommender models like that in LLMs. By tokenizing scenario and task priors into a unified latent space, we transform them from mere auxiliary inputs into active participants within the core feature-interaction backbone. Instead of restricting scenario/task information to the ”shallows” of the model (e.g., gates or heads), tokenized priors can penetrate deep and multi-layered interaction modules. This allows distribution-specific signals to adaptively modulate the model’s massive parameter space, fully harvesting the benefits of the scaling law (corresponding to the first issue). By projecting both input-side scenario contexts and output-side task objectives into a consistent tokenized format, we bridge the gap between MSL and MTL. This enables a holistic and joint modeling framework where scenario-task dependencies are captured through a unified interaction protocol (corresponding to the second issue).
@@ 3.2.2. Feature-Scenario/Task Token Interaction
The interaction between scenario/task tokens and feature tokens primarily aims to enable different scenario tokens or task tokens to utilize distinct feature information, thereby achieving differentiated modeling in multi-scenario and multi-task settings. Since task and scenario information are tokenized from the bottom layer, they can fully interact with feature information layer by layer and from the bottom up. This approach alleviates the challenge of insufficient utilization of complex feature interaction modules, which was previously limited to the certain intermediate or output layers.
Specifically, we define a type of interaction called Domain-aware Attention between scenario/task tokens and feature tokens. Without loss of generality, we take the interaction between task tokens $\bm{T}_{t}$ and feature tokens $\bm{T}_{f}$ as an example to illustrate this module. The Domain-aware Attention can be viewed as a variant of cross-attention. Given the task token $\bm{T}_{t}$, we use it as query, while treating the feature token $\bm{T}_{f}$ as key and value. Then they are interacted through cross multi-head attention as follows:
@@ 3.2.3. Scenario-Task Token Interaction
Specifically, the Domain-fused Module consists of two steps: (1) instance-level scenario token selection, and (2) hybrid information fusion between task and scenario tokens. Firstly, for each instance in the dataset, we select the corresponding scenario tokens $\{{\bm{t}_{s,i},\bm{t}_{s,j}}\}$ from the full scenario token set $\bm{T}_{s}=\{\bm{t}_{s,1},\bm{t}_{s,2},\cdots,\bm{t}_{s,{n_{s}}},\bm{t}_{s,global}\}$ based on the scenario information associated with each instance, along with the global scenario token $\bm{t}_{s,global}$. For example, suppose a model is designed to serve $N_{s}$ scenarios. For a given instance, it must belong to one or more of these scenarios (in cases where scenario overlap exists). Based on this information, the corresponding scenario tokens are selected. After selecting the scenario tokens $\{{\bm{t}_{s,i},\bm{t}_{s,j}},\bm{t}_{s,global}\}$, we aggregate their information using a mean pooling approach:
@@ 3.3. MDL Block
Forward Propagation of Scenario Tokens. The information propagation of scenario tokens primarily involves domain-aware attention between scenario tokens and feature tokens (Eq 8), per-scenario FFN, and residual connection. It can be formulated as follows:
@@
Forward Propagation of Task Tokens. The information propagation of task tokens primarily consists of domain-aware attention between task tokens and feature tokens (Eq 7), domain-fused module (Eq 9, Eq 10), per-task FFN and residual connection. It can be formulated as follows:
@@ 4.3. Ablation Study
The experimental results of the MDL variants are reported in Table 2. From it, we can observe that removing each of the components leads to the performance decrease. The variant $w/o$ task token and $w/o$ scenario token indicate the importance of tokenization of scenario information and task information. The variant $w/o$ global scenario token shows the importance of the global token. The variants $w/o$ task-feautre interaction and $w/o$ scenario-feature interaction indicate the effectiveness of the elaborated interaction mechanisms among these tokens.
@@ 4.5. Online Experiments
To further verify the effectiveness of the proposed approach MDL, we conducted an online A/B test on Douyin Search for one month. The online base model is based on RankMixer (Zhu et al., 2025), with multi-scenario and multi-task learning implemented via MMoE (Ma et al., 2018). The core online metrics include change query rate and LT30, as introduced in Section 4.1.2. Table 3 reports the relative improvements in the different scenarios and the overall performance. The proposed method yielded +0.0626% improvement in LT30 and -0.3267% reduction in change query rate. As for each scenario, it also achieved significant improvements on the core metrics.
@@ 5. Related Work
Large-scale Recommendations. Inspired by the success of scaling laws in large language models (LLMs) (Hoffmann et al., 2022; Kaplan et al., 2020), the scaling laws of recommendation models have recently been widely studied and have achieved significant success in real-world industrial recommendation systems (Zhang et al., 2024a; Zhu et al., 2025; Yan et al., 2025; Zhai et al., 2024; Han et al., 2025a; Zhang et al., 2025; Huang et al., 2026). Typically, these methods increase the model parameter size and computational complexity by stacking various efficient feature interaction structures, thereby improving the recommendation performance. One type of method focuses on scaling the parameters and complexity of the non-sequential feature interaction module for better performance. For example, Wukong (Zhang et al., 2024a) scales the feature interaction module by introducing a stacked factorization machine-based architecture. Rankmixer (Zhu et al., 2025) introduces multi-head token mixing and pertoken feed-forward networks to scale recommender models, while maintaining inference latency and achieving superior performance. FAT (Yan et al., 2025) scales recommendation models by introducing field-aware transformer with field-based interaction priors in attention. Another typical method adopts the transformer-based structure for modeling non-sequential and sequential feature in a unified way and scaling constantly. For instance, HSTU (Zhai et al., 2024) sequentializes the heterogeneous feature space and then performs sequence transduction conditioned on contextual and candidate signals. MTGR (Han et al., 2025a) further scales recommendation models based on HSTU for acceleration and leakage avoidance. OneTrans (Zhang et al., 2025) introduces a unified transformer-based backbone that simultaneously models user-behavior sequences and feature interactions via a unified tokenizer, enabling efficient parameter scaling and superior performance. Compared to these methods, our work further explores how large-scale recommendation models can better perform in multi-scenario and multi-task settings.
@@
Multi-scenario and Multi-task Recommendations. In real-world industrial recommendation systems, a recommendation model usually needs to serve multiple scenarios and tasks. Multi-scenario learning (MSL) and multi-task learning (MTL) in recommender system have been widely studied in previous works (Caruana, 1998; Ma et al., 2018; Misra et al., 2016; Ma et al., 2019; Tang et al., 2020; Li et al., 2020; Sheng et al., 2021; Chang et al., 2023; Zhou et al., 2023; Song et al., 2024; Yan et al., 2024; Ding et al., 2021). For MTL, most existing methods adopt a shared-specific architecture in which task-shared and task-specific components are combined together (e.g., Sharedbottom (Caruana, 1998), Cross-stitch (Misra et al., 2016), SNR (Ma et al., 2019)). In addition, Mixture-of-Experts (MoE) based structures have also been widely proposed, aiming to enhance the effectiveness of multi-task learning by designing various expert mechanisms (e.g., MMoE (Ma et al., 2018), PLE (Tang et al., 2020)). For MSL, it aims to use a unified model to learn data from multiple scenarios, thereby improving recommendation performance. Most existing work also adopts a shared-specific architecture in which scenario-shared and scenario-specific components are combined together. The framework used in MTL can also be applied in MSL by treating the predictions for different scenarios as distinct prediction tasks (e.g., SharedBottom (Caruana, 1998), MMoE (Ma et al., 2018)). In addition, some specialized multi-scenario methods have also been proposed (Li et al., 2020; Sheng et al., 2021; Chang et al., 2023; Zhou et al., 2023; Song et al., 2024; Mu et al., 2023; Zhang et al., 2022a; Yang et al., 2022; Zhang et al., 2022b; Niu et al., 2021; Shen et al., 2021; Chen et al., 2020). For example, HMoE (Li et al., 2020) proposes the hybrid of implicit and explicit MoE structure which learns the scenario and task relationships. STAR (Sheng et al., 2021) introduces a start topology structure where a centered network is used for all scenarios and scenario-specific network is generated for each scenario. PEPNet (Chang et al., 2023) takes the scenario related prior information to adjust the feature embeddings and hidden units through gate mechanisms. In contrast to these approaches which are based on the shared-specific architecture, we propose a novel multi-distribution learning framework from the perspective of tokenization, aiming to enhance the learning capability and unified modeling ability of large-scale recommendation models in multi-scenario and multi-task settings.
@@ 6. Conclusion
In this paper, we propose a novel multi-distribution learning framework for multi-scenario learning (MSL) and multi-task learning (MTL) in large-scale industrial recommender system, named MDL. Unlike existing MSL and MTL recommendation methods that are based on a shared-specific architecture, the proposed approach inspired by large language models (LLMs) is to leverage prompt-based tokenization to achieve model predictions in various scenarios and tasks settings. Specifically, MDL tokenizes feature, scenario, and task information into a unified tokenized format, and elobrates three key interaction mechanisms among these tokens. Extensive experiments on both offline evaluation and online test have demonstrated the effectiveness. The proposed approach MDL has been fully deployed on Douyin Search platform.